OpenStudy (anonymous):

An old house in Boxerville measures 30 feet in height and has windows 5-feet in width. An architecture student builds a proportional model that is 3-feet high. How wide are the model's windows, in feet?

OpenStudy (triciaal):

ratio similar figures compare window/height let the value of the model's window = x 5/30 = x /3 solve for x model's window = 1 ft